# Identifying companies and financial actors exposed to marine tipping points

### v1

@@ Discussion
Recent studies showed that heatwaves alone are unlikely to decrease fish biomass36 despite some events have been already observed8,10,35. Yet, it is often the effect of compound events or the interaction of multiple drivers that can drive abrupt fish declines or other marine regime shifts2,13. Their frequency and intensity are predicted to increase in the near future with climate change scenarios1,8,9,13. In this paper, we examined which corporate and financial actors are exposed to marine tipping points by identifying vessels and companies operating in areas that are simultaneously exposed to marine heat waves and instabilities in primary productivity.
@@
Here we reported lists of top countries, companies, and their shareholders associated with vessels fishing in areas at high risk of marine regime shifts. In summary, we found that fishing fleets associated with China, the US, Japan and Norway are particularly exposed. The shareholders of companies exposed to tipping point risks are based in Norway, China, Spain, Russia, the UK and Denmark. Spain, UK and France are home to particularly vulnerable shareholders investing in companies who fish in high-risk areas, while Russia is headquarters for fishing companies getting such investments. We developed a method to track vessels and companies at risk which depends on a mix of publicly available43 and non-public datasets. Our method and findings are limited by data completeness and transparency.
@@
While transparency and accountability are values sought after by the financial sector, our analysis highlights the complexity and opacity of global fisheries, even with access to state-of-the-art data42,43 and databases (e.g. Lloyds, Orbis). We first identified close to 17,000 vessels fishing in high-risk areas, but as we traced ownership up the chain –from vessel owner to company beneficiaries and shareholder– data availability dropped by an order of magnitude at each step. A high proportion of vessels have unidentified flags (20%) and lack registered owners (81.41%). Thus our results are biassed towards the vessels and countries that report ownership data, that is primarily publicly listed companies operating in legal fisheries. Further research is needed to correct this bias45. Nonetheless, our study serves as a benchmark for what is achievable with current databases and is the first study to connect ocean tipping points to financial risks.
@@
Another limitation stems from known problems with the AIS data on which our vessel identification is based. The data is known to be less reliable before 2017 and not all countries follow the same standards in reporting. For example, the European Union mandates reporting for all vessels over 15 metres long, while China mandates the usage of AIS in all their vessels. Additionally, AIS is primarily used by vessels over 300 gross tons which excludes many small-scale fisheries vessels and some commercial ones. Although these limitations have led previous studies to restrict their analyses to the high seas43, we included exclusive economic zones (EEZs) in our analysis because they are often prone to marine heatwaves and regime shifts (Fig 1). In fact, over 500 observed hypoxic and anoxic zones –a regime shift that impacts primary productivity and hence fishing– have been observed in national jurisdictions6,7.
@@
Despite these limitations, our study offers key insights for managing the financial risks associated with marine tipping points. In contrast to previous work that identified keystone actors dominating the exploitation of natural resources16,26,27, we found that the fisheries sector does not seem to be highly concentrated, at least not at the level of decision-making that is exposed to risk of regime shifts – where to go fishing and when. Market concentration probably occurs further down the supply chain, at the level of intermediaries29, but this level of aggregation makes it harder for financial institutions and regulators to discriminate whether the fish traded comes from areas prone to tipping points.
@@
Financial actors and corporations are increasingly required to report their climate and nature-related risks, as seen in initiatives like the Task Force on Climate-Related Disclosures (TCFD), or the European Corporate Reporting Directive (CSRD). Central to these reporting efforts is the concept of double materiality: companies need to assess what is their vulnerability to ecological or climate change; but also how their activities impact climate change or ecological degradation46. This feedback loop heightens the risks faced by both the company and its investors. Our maps of tipping point risk combined with readily available georeferenced information on fishing fleet location enable companies and financial investors to assess their exposure, adjust their fishing areas (companies), and manage their investment portfolio (investors) to reduce exposure. Heatwaves and cold spells forecasting tools9, as well as near-real-time monitoring of ecosystem resilience38 can support adaptive strategies that minimize risk exposure.
@@
When fisheries collapse (e.g. cod in the 1990s) or when stocks decline abruptly (e.g. the Blob in 2014-19), governments often need to step up and pay the economic and social costs. At the same time, governments are subsidizing high seas fisheries, without such subsidies 54% of their catch would have been unprofitable otherwise14. These subsidies need to stop or redirected to economic activities without impacts on ecological integrity and biodiversity23. But besides the already-known role of government subsidies14,23, our study shows that governments also have a role to play as shareholders in some of the biggest fishing fleets. This is particularly evident in China where national, regional and municipal government organizations sit on the boards of fishing companies exposed to tipping point risks. These governmental and public organizations have decision power to mitigate risk by directing fisheries to less risky grounds or by holding the fleet accountable for avoiding high-risk areas.
@@
As heatwaves and marine regime shifts unfold, there is an additional risk that governance frameworks may become outdated. While benthic species tend to have low mobility and be impacted by these events, pelagic species can more readily relocate if the area affected is within their mobility range. The problem arises when a stock quota that was assigned in an economic zone to a particular set of actors moves to another economic jurisdiction, generating conflicts between fisheries and their managing institutions47,48. Many fish stocks are already expected to move polewards due to climate change49, and marine regime shifts may exacerbate these tensions. Governments and fishing authorities should put mechanisms in place to prevent, mediate or resolve these emerging conflicts, as non-linear environmental changes can not only alter the “rules of the game” but also shift the playing field.
@@ Conclusion
Many environmental problems stem from how governments, cities and businesses exploit natural resources. We show that a variety of these actors, and their financiers, are themselves exposed to the risk of marine tipping points, which threaten the productivity of marine ecosystems. Redesigning subsidy programs, lobbying company boards to fish in low-risk areas, or supporting the implementation of international agreements that take into account potential range shifts and conflicts are just a few examples of how these actors can incentivise practices that are aligned with long term sustainability values.
@@ Acknowledgements
This work has been supported by the project Networks of Financial Rupture funded by FORMAS grant 2020-00198; and the Finance to Revive Biodiversity (FinBio) program, financed by Mistra – the Swedish Foundation for Strategic Environmental Research (DIA 2020/10). Our work benefited from discussions with Robert Blasiak.
@@
1 null -700.78 1403.55
2 difference -681.59 1379.18
3 full -654 1336

### v2

@@ Discussion
Recent studies showed that heatwaves alone are unlikely to decrease fish biomass40 despite some events have been already observed8,10,39. Yet, it is often the effect of compound events or the interaction of multiple drivers that can drive abrupt fish declines or other marine regime shifts2,13. Their frequency and intensity are predicted to increase in the near future with climate change scenarios1,8,9,13. In this paper, we examined which corporate and financial actors are exposed to marine tipping points by identifying vessels and companies operating in areas that are simultaneously exposed to marine heat waves and instabilities in primary productivity. We focused on heat waves and changes in primary productivity because of data availability, acknowledging that marine ecosystems can tip due to other causes as well2.
@@
Here we reported lists of top countries, companies, and their shareholders associated with vessels fishing in areas at high risk of marine regime shifts. In summary, we found that fishing fleets associated with China, the US, Japan and Norway are particularly exposed. The shareholders of companies exposed to tipping point risks are based in Norway, China, Spain, Russia, the UK and Denmark. Spain, UK and France are home to particularly vulnerable shareholders investing in companies who fish in high-risk areas, while Russia is headquarters for fishing companies getting such investments. The implications are that the management is not only in the hands and decisions of companies. Countries could also design laws and regulations that restrain or disincentivise investments in other regions of the world where fishing activities are occurring in risk areas. We developed a method to track vessels and companies at risk which depends on a mix of publicly available55 and non-public datasets.
Our method and findings are limited by data completeness and transparency. For example, corporate ownership often involves shell companies, offshore entities and indirect holdings that may not be fully documented in Orbis, or risk misidentifying or omitting ultimate owners. Another example is vessel flags, many vessels operate with flags of convenience obscuring the true ownership structure in the data. The absence of financial data at vessel level further weakens the link between company ownership and fishing activities.
@@
While transparency and accountability are values sought after by the financial sector, our analysis highlights the complexity and opacity of global fisheries, even with access to state-of-the-art data54,55 and databases (e.g. Lloyds, Orbis). We first identified close to 17,000 vessels fishing in high-risk areas, but as we traced ownership up the chain –from vessel owner to company beneficiaries and shareholder– data availability dropped by an order of magnitude at each step. A high proportion of vessels have unidentified flags (20%) and lack registered owners (81.41%). Thus our results are biased towards the vessels and countries that report ownership data, that is primarily publicly listed companies operating in legal fisheries. Further research is needed to correct this bias60. Nonetheless, our study serves as a benchmark for what is achievable with current databases and is the first study to connect ocean tipping points to financial risks.
@@
Another limitation stems from known problems with the AIS data on which our vessel identification is based. The data is known to be less reliable before 2017 and not all countries follow the same standards in reporting. For example, while AIS is primarily used by vessels over 300 gross tons operating beyond national jurisdiction, a country like China mandates AIS for a broad range of vessels also within the Exclusive Economic Zone (EEZ). This has led previous studies to restrict their analyses to the high seas55. Here, we nevertheless included EEZs in our analysis because they are often prone to marine heatwaves and regime shifts (Fig 1). In fact, over 500 observed hypoxic and anoxic zones –a regime shift that impacts primary productivity and hence fishing– have been observed in national jurisdictions6,7.
@@
Despite these limitations, our study offers key insights for managing the financial risks associated with marine tipping points. In contrast to previous work that identified keystone actors dominating the exploitation of natural resources17,30,31, we found that the fisheries sector does not seem to be highly concentrated, at least not at the level of decision-making that is exposed to risk of regime shifts – where to go fishing and when. Market concentration probably occurs further down the supply chain, at the level of intermediaries33, but this level of aggregation makes it harder for financial institutions and regulators to discriminate whether the fish traded comes from areas prone to tipping points.
@@
Financial actors and corporations are increasingly required to report their climate and nature-related risks, as seen in initiatives like the Task Force on Climate-Related Disclosures (TCFD), or the European Corporate Reporting Directive (CSRD). Central to these reporting efforts is the concept of double materiality: companies need to assess what is their vulnerability to ecological or climate change; but also how their activities impact climate change or ecological degradation61. This feedback loop heightens the risks faced by both the company and its investors. Our maps of tipping point risk combined with readily available georeferenced information on fishing fleet location enable companies and financial investors to assess their exposure, adjust their fishing areas (companies), and manage their investment portfolio (investors) to reduce exposure. Heatwaves and cold spells forecasting tools9, as well as near-real-time monitoring of ecosystem resilience42 can support adaptive strategies that minimize risk exposure.
@@
When fisheries collapse (e.g. cod in the 1990s) or when stocks decline abruptly (e.g. the Blob in 2014-19), governments often need to step up and pay the economic and social costs. At the same time, governments are subsidizing high seas fisheries, without such subsidies 54% of their catch would have been unprofitable otherwise15. Our study shows that governments also have a role to play as shareholders in some of the biggest fishing fleets. This is particularly evident in China where national, regional and municipal government organizations sit on the boards of fishing companies exposed to tipping point risks. These governmental and public organizations have decision power to mitigate risk by directing fisheries to less risky grounds or by holding the fleet accountable for avoiding high-risk areas.
@@
As heatwaves and marine regime shifts unfold, there is an additional risk that governance frameworks may become outdated. While benthic species tend to have low mobility and be impacted by these events, pelagic species can more readily relocate if the area affected is within their mobility range. The problem arises when a stock quota that was assigned in an economic zone to a particular set of actors moves to another economic jurisdiction, generating conflicts between fisheries and their managing institutions62,63. Many fish stocks are already expected to move polewards due to climate change64, and marine regime shifts may exacerbate these tensions. Governments and fishing authorities should put mechanisms in place to prevent, mediate or resolve these emerging conflicts, as non-linear environmental changes can not only alter the “rules of the game” but also shift the playing field.
@@ Conclusion
In this paper, we asked who are these corporate and financial actors exposed to marine tipping points? We first identified areas of the ocean where tipping points can manifest because they are already showing a high intensity of extreme and severe heat wave events, or they are already showing statistical signatures of resilience loss in primary productivity. We found that about 30% of the oceans are at risk of regime shifts (extreme marine heat waves and signals of instability in primary productivity), and provided evidence of past regime shifts from literature reviews, and abrupt change observed in primary productivity and stock assessments data. 16% of the oceans have shown abrupt declines in primary productivity, 42% if accounting only the linear negative trend; while 252 fish stocks resulted in abrupt declines on historical data. The literature already showcase past events with impacts on employment and economic losses on the millions of dollars. So who is at risk today?
We found that 15% (16,878) of the vessels tracked in Global Fishing Watch do actually fish in areas susceptible to marine regime shifts. We did not find the levels of market concentration found in other economic sectors such as aquaculture, in fact the ownership structure is quite sparse with most companies owned by one or two shareholders. Nonetheless, we find that Norway and China are important actors in hosting many of the vessels and shareholders in our data, but they are not strongly involved on foreign investment. Spain, the UK and France are central countries of origin for foreign investment, while Russia is a hotspot attracting these investments. The likelihood of these financial flows seems to be related to differences in government effectiveness and political stability.
We provided lists of top shareholders and top companies with vessel ownership fishing in areas prone to marine regime shifts. We discussed the limitations of our study mainly related to data transparency. Nonetheless, we hope these lists of companies and shareholders help to identify actors with agency to manage the risk of marine regime shifts. We discussed potential interventions that can reduce regime shifts risks. Redesigning subsidy programs, lobbying company boards to fish in low-risk areas, or supporting the implementation of international agreements that take into account potential range shifts and conflicts are just a few examples of how these actors can incentivise practices that are aligned with long term sustainability values.
@@ Acknowledgements
This work has been supported by the project Networks of Financial Rupture funded by FORMAS grant 2020-00198; and the Finance to Revive Biodiversity (FinBio) program, financed by Mistra – the Swedish Foundation for Strategic Environmental Research (DIA 2020/10), and the Swedish Research Council (VR grant 2022-04122). Our work benefited from discussions with Robert Blasiak.
@@
Figure S1: Empirical evidence of regime shifts. Over 1440 marine regime shifts are documented on the regime shifts database, 788 cases are of hypoxia, followed by marine eutrophication (245) and fisheries collapse (181). For each dot in the map there is a paper in the scientific literature reporting regime shifts in marine ecosystems. Source: www.regimeshifts.org (beta version at www.regimeshifts.netlify.app).
Figure S2: Linear changes in primary productivity. We estimated Kendall tau on the full time series of Chlorophyll-A at 9Km resolution. 77.3% of the ocean shows a slope significantly different from zero, 41.9% with a declining trend (A). In (B) the same time series have been detrended using seasonal-trend decomposition using LOESS, and a linear model fit on the residuals showing that there are still negative and significant trends in the central Pacific and Indian oceans that are not related to seasonality.
Figure S3: Linear changes in fish stocks. We first compared the timing of the break points detected with the Pettitt test and the structural change test (A). Despite both test agree on the existence of the shift, there are some differences in timing. The median decline is 50% (red dotted line in B-C), and most shifts occurred between the 1970s and early 2000s. A comparison with the Sen slope test show that linear trends agree with the direction of change reported on the abrupt changes, the units of the slope are in biomass.
@@
1 null -537.22 1076.43
2 difference -527.78 1073.56
3 full -516.75 1065.5